One third of a number is 4 more than 11. What is the number?!
Answer by Theo(13342)

let x be the number.
1/3 * x = 11 + 4 = 15
solve for x to get x = 3 * 15 = 45.
that's your number.
1/3 of 45 = 15.
15 is 4 more than 11.